OpenStudy (poofypenguin):

Solve the following inequality. x(x−7)/[(x^2)−10x−39] <= 0

OpenStudy (poofypenguin):

Here's the inequality in proper notation: \[\frac{ x(x-7) }{ x^{2}-10x-39 }\le0\]

OpenStudy (anonymous):

if you multiply both sides by the denominator, all you have is x(x-7) <= 0

OpenStudy (poofypenguin):

i tried that, but it didn't work out to the write answer...

OpenStudy (anonymous):

is x <=7 an answer? because x(x-7) <= 0, then x <=0 and x <=7 the latter includes the former so it's just x <=7

OpenStudy (poofypenguin):

i'm not actually sure what the answer is... i'm typing it into an online worksheet and it's just telling me that that's the answer.

Directrix (directrix):

-3 < x ≤ 0, 7 ≤ x < 13 @PoofyPenguin